OpenStudy (anonymous):

Select all of the following that are ordered pairs of the given function. f(x) = 3 - 2x (-2, -1) (-1, 5) (0, 3) (1, 0) (2, -1)

OpenStudy (anonymous):

f(x) = 3 - 2x or 3-2(x) = y which of these are true? 3-2(-2) = -1 T/F? 3-2(-1) = 5 T/F? 3-2(0) = 3 T/F? 3-2(1) = 0 T/F? 3-2(2) = -1 T/F?